# SC-Diff: 3D Shape Completion with Latent Diffusion Models

### v1

@@ Abstract
This paper introduces a 3D shape completion approach using a 3D latent diffusion model optimized for completing shapes, represented as Truncated Signed Distance Functions (TSDFs), from partial 3D scans. Our method combines image-based conditioning through cross-attention and spatial conditioning through the integration of 3D features from captured partial scans. This dual guidance enables high-fidelity, realistic shape completions at superior resolutions. At the core of our approach is the compression of 3D data into a low-dimensional latent space using an auto-encoder inspired by 2D latent diffusion models. This compression facilitates the processing of higher-resolution shapes and allows us to apply our model across multiple object classes, a significant improvement over other existing diffusion-based shape completion methods, which often require a separate diffusion model for each class. We validated our approach against two common benchmarks in the field of shape completion, demonstrating competitive performance in terms of accuracy and realism and performing on par with state-of-the-art methods despite operating at a higher resolution with a single model for all object classes. We present a comprehensive evaluation of our model, showcasing its efficacy in handling diverse shape completion challenges, even on unseen object classes. The code will be released upon acceptance. ${}^{\ast}$ Xingxing Zuo is the corresponding author.
Keywords:
Object Shape Completion Diffusion Model TSDF
Figure 1: Our method produces realistic and accurate object shape completions from realistic partial scans with a higher degree of detail than state-of-the-art methods, including 3D-EPN [12] and AutoSDF [22]. Image input is optional in our work.
@@ 1 Introduction
The evolution of digital technologies has revolutionized the role of 3D content, particularly in areas like augmented/virtual reality, gaming, robot navigation, and industrial design. These domains rely heavily on the creation and manipulation of 3D models, a process that remains technically challenging and labor-intensive, often requiring specialized skills and tools.
A significant development in this landscape in recent years has been the widespread of commodity RGB-D sensors (e.g., Intel RealSense, Microsoft Kinect, iPhone, etc.). These sensors have enabled remarkable progress in 3D reconstruction and understanding, facilitating the generation of detailed 3D models from the real world. The ability to use these sensors for high-quality scanning has far-reaching implications, from enhancing virtual reality experiences to improving the accuracy of robot navigation and streamlining design processes in various industries.
However, despite advancements in reconstruction quality [1, 25], the scans obtained from RGB-D sensors are often partial or fragmented, leaving gaps in reconstructing complete 3D models. This fundamental challenge arises due to environmental obstacles and the restricted sensor reach, both of which impede the attainment of complete scanning. Addressing this challenge is vital for fully utilizing 3D scanning technologies and maximizing the potential of 3D content in the aforementioned fields.
Recently, a variety of learning-based methods for 3D shape completion have emerged, utilizing techniques such as fully convolutional networks [13] and transformer based architectures [38]. These approaches typically focus on establishing a direct mapping function between partial scans and their complete counterparts. However, they can sometimes encounter limitations, such as overfitting and the potential introduction of artifacts. In contrast, probabilistic methods approach shape completion as a conditional generation task, offering the flexibility to produce multiple plausible outcomes. This category includes innovative solutions like variational auto-encoders [22], adversarial training [41], and diffusion models [6, 23, 8, 7], which collectively represent a significant advancement in generating diverse and realistic shape completions in scenarios with inherent ambiguity.
Diffusion models have achieved significant success in various image generation tasks, especially in text-conditional and image-conditional generation [15, 31]. Advances such as ControlNet [42] have introduced spatially consistent conditioning mechanisms, enhancing control over the spatial composition of generated images. This makes diffusion models highly promising for the conditional completion of 3D shapes. Latest methods like SDFusion [6] and DiffRF [6] conceptualize shape completion as filling artificially generated missing cubic regions, but this does not fully address the challenges of realistic sensor-captured 3D shapes, which often have varying levels of incompleteness and noise [6, 23].
In this paper, we present an approach for 3D shape completion from partial scans, aiming to produce realistic, high-fidelity, complete shapes. We complete a Truncated Signed Distance Function (TSDF) based 3D shape representation with a latent diffusion model, integrating image-based and spatially consistent conditioning inspired by ControlNet [42]. Our approach utilizes an auto-encoder architecture trained on both 3D and 2D objectives [36, 14] to compress the input TSDF into a lower dimensional latent space, enabling the processing of higher resolution shapes efficiently in the diffusion model (moving from $32^{3}$ to $64^{3}$ TSDF voxel grids). When compared to a non-latent approach, our latent-diffusion model uses $30.72\%$ less GPU Memory. This addresses the common challenge of high computational demands that limit existing 3D generative models to lower resolutions [8, 6].
The key contributions of this work are:
  • Compression of 3D shapes into a latent space, enabling the processing of voxel grids of higher resolutions and the learning of diffusion-based shape completion for multiple classes with a single model.
  • Involving both 3D supervision and volume rendering-enabled 2D supervision for the learning of the Vector Quantised Variational AutoEncoder (VQ-VAE) resulting in a compact representation of the input TSDFs.
  • Incorporating two independent and complementary conditioning mechanisms: image-based conditioning with cross-attention and spatial conditioning, integrating 3D features from partial scans inspired by DiffComplete  [8].
@@ 2 Related Work
Traditional approaches to 3D shape completion primarily focus on repairing minor imperfections and geometrical gaps, employing strategies like Laplacian hole filling [17] and Poisson surface reconstruction [4]. While effective for filling small gaps in the 3D shapes, these methods are of limited use in scenarios with a larger scale of gaps in the input scans. However, the advent of extensive 3D datasets has led to the development of learning-based approaches. While retrieval-based methods [3, 40, 2] extract the most fitting shapes from a database to complete the incomplete input, learning-based fitting methods [39, 12, 10, 14, 37] that aim to minimize the disparity between network-predicted shapes and actual ground truth have become more prevalent. Examples of such methods are 3D-EPN [12] and Scan2Mesh [11] which have employed 3D encoder-decoder architectures for predicting complete shapes from partial volumetric data. More recently, PatchComplete [29] introduces a novel design by learning multi-resolution patch priors, enhancing the completion of shapes, particularly for previously unseen categories. While these models can provide high-accuracy completions, their deterministic nature may lead to overfitting. Furthermore, they often lack the ability to generate finer details in shapes.
@@
Generative methods, including Generative Adversarial Networks (GANs) [41] and AutoEncoders [22], have recently been introduced to the field. These methods excel in generating a diverse range of plausible shapes from partial inputs. However, this often comes at the cost of accuracy, especially when a precise ground truth is available. Diffusion models have emerged as a popular family of generative models, renowned for their impressive quality, diversity, and expressiveness in tasks such as image synthesis [31], image editing [20], and text-to-image synthesis [28]. While extensively researched for 2D data, their application to 3D data has only been explored very recently, using different 3D representations such as point-clouds [19, 44] neural radiance fields [21], or signed distance functions [7, 18]. SDFusion [6], specifically, introduces a latent diffusion model for shape generation aimed at mitigating the computational challenges associated with 3D representations, and conditions it on various modalities like images, text, and shapes. However, SDFusion merely regards completion as extending a partial but complete shape by filling artificially cropped out 3D boxes. In contrast, our approach does not make any assumption on the partial shape as real-world scans are characterized by noise and varying degrees of incompleteness. Another approach, DiffComplete [8], offers a more realistic shape completion method but directly works on the 3D space and is therefore restricted to lower-resolution voxel grids of size $32^{3}$. As opposed to DiffComplete, our approach can achieve much higher resolution by encoding the 3D space to compact latent codes. Furthermore, most previous methods [29, 23, 21] work on the assumption that the object class is known prior to the completion, which makes them hard to deploy in real-world applications. In contrast, our model is class-agnostic, and outputs high-detailed, realistic completions, even for unseen classes.
@@ 3 Methodology
Figure 2: We train a Vector Quantized Variational Autoencoder (VQ-VAE) to encode complete TSDF voxel grids $\mathbf{X}$ into a compact latent variable $\mathbf{z}$. Then, we use this learned latent space to train our conditional shape completion diffusion model. First, we corrupt the clean latent code z through a forward diffusion process $q(\mathbf{z}_{t}|\mathbf{z}_{-1})$ by gradually adding Gaussian noise until reaching the normally distributed variable $\mathbf{z}_{T}$. Our diffusion model is then trained to learn the reverse diffusion process $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})$ conditioned on a partial TSDF scan c of the shape by aggregating features from a control branch and/or a partial RGB image $\mathbf{I}$ of the shape by applying cross-attention on CLIP Features of the given images. This process allows us to recover the latent code $\mathbf{\hat{z}}$, and finally, using the VQ-VAE’s decoder, the completed shape $\mathbf{\hat{X}}$.
We approach 3D shape completion as a generative task based on a diffusion probabilistic model, aiming to produce a complete shape $\hat{\mathbf{X}}$ from a given partial scan represented by an RGB Image $\mathbf{I}$ and a partial 3D TSDF $\mathbf{c}$. While diffusion models have proven highly effective in generating high-resolution images, their direct application to high-resolution 3D shapes is hindered by the substantial demands on computation and memory. Consequently, 3D diffusion-based approaches are often confined to low-resolution voxel grids, typically around $32^{3}$  [6, 22]. To address this limitation, we first compress the 3D shape into a compact latent space using a Vector Quantized Variational Autoencoder (VQ-VAE), as detailed in Sec. 3.1. This VQ-VAE is trained with supervision from both 3D TSDF, and 2D depth and normal renderings from a raycasting-based volume rendering approach [14]. This compression strategy allows us to implement our diffusion model in a lower-dimensional space, a process referred to as Latent Diffusion (Sec. 3.2). Our model includes two distinct conditioning mechanisms (Sec. 3.3) in the generative process: the first leverages RGB Image $\mathbf{I}$’s CLIP features  [27] via cross attention, and the second uses spatially localized conditions [42, 8] derived from the 3D partial scans $\mathbf{c}$. An overview of our approach is provided in Fig. 2.
@@
3.1.1 Vector Quantized Variational Autoencoder
We employ a 3D Variant of the Vector Quantized Variational Autoencoder (VQ-VAE)  [36, 30] for condensing the compressed latent spaces of various shapes. The architecture of VQ-VAE consists of an encoder $\mathcal{E}$, a decoder $\mathcal{D}$, and a codebook $\bm{\mathcal{Z}}\in\mathbb{R}^{K_{\mathcal{Z}}\times D}$, with $K_{\mathcal{Z}}$ and $D$ representing the codebook’s length and the dimensionality of each latent vector, respectively.
For a given TSDF volume $\mathbf{X}$, encoding into a latent representation is done as $\mathbf{z}=\mathcal{E}(\mathbf{X})$. The decoder then reconstructs $\mathbf{X^{\prime}}=\mathcal{D}(Q(\mathbf{z}))$ from $\mathbf{z}$, where $Q(\cdot)$ represents the quantization step, mapping each of the latent vectors $\mathbf{z}_{i}$ in $\mathbf{z}$ to the nearest codebook vector $\bm{\mathcal{Z}}_{j}$ using $Q(\mathbf{z}_{i})=\bm{\mathcal{Z}}_{j}$, where $j=\arg\min_{k\in[1,K_{z}]}||\mathbf{z}_{i}-\bm{\mathcal{Z}}_{k}||_{2}$. Here, $\bm{\mathcal{Z}}_{k}$ corresponds to $k$th elements of $\bm{\mathcal{Z}}$.
We pre-train our VQ-VAE loss with the following 3D losses:
  • Reconstruction Loss minimizes the difference between the input and reconstructed TSDF voxel grid: $\mathcal{L}_{R}=||\mathbf{X}-\mathbf{X^{\prime}}||_{1}$
@@ 3.1.1 Vector Quantized Variational Autoencoder
  • Codebook Loss ensures the chosen codebook vector aligns with the encoder output:
@@
Alongside 3D losses, our methodology incorporates 2D supervision from rendering the reconstructed TSDF grids $\mathbf{X^{\prime}}$, inspired by SPSG [14]. We render both depth ($\mathbf{D_{v}^{\prime}}$) and normal images ($\mathbf{N_{v}^{\prime}}$) from $\mathbf{X^{\prime}}$ at each camera view $v$ in a differentiable manner via raycasting-based volume rendering.
The 2D supervision losses employed for VQ-VAE training include:
  • 2D Reconstruction Loss: An $l_{1}$ loss guiding depth and normal reconstruction:
    $\mathcal{L}_{R_{2D}}=||\mathbf{D_{v}}-\mathbf{D_{v}^{\prime}}||_{1}+||\mathbf{N_{v}}-\mathbf{N_{v}^{\prime}}||_{1}$ (1)
    where $\mathbf{D_{v}}$ and $\mathbf{N_{v}}$ correspond to the ground-truth depth and normal values, and $\mathbf{D_{v}^{\prime}}$ and $\mathbf{N_{v}^{\prime}}$ are the rendered depth and normal images.
  • 2D Adversarial Loss: A patch-based adversarial loss on normals $\mathbf{N_{v}^{\prime}}$ for realistic geometry reconstruction:
@@
    where $\mathrm{Disc}(\cdot)$ denotes the CNN based discriminator network.
The final VQ-VAE objective combines both 3D and 2D losses:
$\mathcal{L}_{VQVAE}=\mathcal{L}_{R}+\beta\mathcal{L}_{C}+(1-\beta)\mathcal{L}_{CB}+\gamma_{R}\mathcal{L}_{R_{2D}}+\gamma_{A}L_{A_{2D}},$ (3)

### v2

@@ Abstract
We present a novel 3D shape completion framework that unifies multimodal conditioning, leveraging both 2D images and 3D partial scans through a latent diffusion model. Shapes are represented as Truncated Signed Distance Functions (TSDFs) and encoded into a discrete latent space jointly supervised by 2D and 3D cues, enabling efficient high-resolution processing while reducing GPU memory usage by 30% compared to state-of-the-art methods. Our approach guides the generation process with flexible multimodal conditioning, ensuring consistent integration of 2D and 3D information from encoding to reconstruction. Our training strategy simulates realistic partial observations, avoiding assumptions about input structure and improving robustness in real-world scenarios. Leveraging our efficient latent space and multimodal conditioning, our model generalizes across object categories, outperforming class-specific models by 12% and class-agnostic models by 47% in $l_{1}$ reconstruction error, while producing more diverse, realistic, and high-fidelity completions than prior approaches.
@@ 1 Introduction
The rapid evolution of digital technologies has made 3D content a cornerstone in augmented/virtual reality, gaming, robotics, and industrial design. These applications rely on complete and accurate 3D models, but creating such models remains challenging and labor-intensive. Commodity RGB-D sensors (e.g., Intel RealSense, Microsoft Kinect, iPhone LiDAR) have enabled significant progress in capturing real-world 3D data. However, despite advancements in reconstruction quality [2, 29], the acquired scans are typically partial, incomplete, or noisy due to occlusions and limited sensor viewpoints, leaving large gaps in the reconstruction pipeline. Addressing this fundamental challenge is essential for many downstream technologies.
Figure 1: Our method, SC-Diff, is a latent diffusion framework for 3D shape completion that leverages 2D and 3D conditioning across the entire pipeline, scales to high-resolution TSDFs, and generalizes across object categories, achieving state-of-the-art accuracy, diversity, and realism.
A central design choice in tackling 3D shape completion is the representation of 3D geometry. Signed Distance Functions (SDFs), meshes, point clouds, and neural radiance fields each offer unique advantages, but SDF-based volumetric representations remain a robust option for high-fidelity completion. Traditional learning-based completion methods [14, 42] often map partial scans directly to full shapes. While effective to some extent, such deterministic mappings tend to overfit and fail to capture the inherent ambiguity of missing geometry. Generative approaches instead model shape completion as a conditional generation problem, yielding more diverse and plausible results [25, 45, 7, 9]. Diffusion models in particular have emerged as powerful probabilistic generative frameworks, offering high-quality and diverse outputs [18, 35].
Despite this progress, existing diffusion-based shape completion methods face several critical limitations. First, they typically adopt simplified training setups such as filling artificially cropped cubic regions [7], which do not reflect realistic sensor-induced incompleteness. Second, they do not scale to high resolutions. Current state-of-the-art methods operate at $32^{3}$ voxel grids [7, 9] since dense latent grids and memory-hungry architectures prohibit scaling to larger volumes. As a result, fine-grained details are lost, and applying these methods to real-world, high-resolution 3D scans remains infeasible. Third, they often operate under restrictive unimodal conditioning [9, 25, 32] or assume multi-view inputs, which largely limits real-world applications, as obtaining sufficient coverage to capture geometric detail is infeasible and comes with extensive computational and memory costs [17]. To overcome these challenges, we propose SC-Diff, a framework for realistic, multimodal, and scalable 3D shape completion. Our goal is to generate high-fidelity, complete shapes while integrating multimodal inputs throughout the pipeline. Specifically, we complete Truncated Signed Distance Function (TSDF) representations using a latent diffusion model conditioned on both image and spatial information. Our contributions are:
  • We are the first to unify latent diffusion with both 2D and 3D conditioning, applied consistently from encoder supervision to the denoising process, for object completion. We introduce a discrete latent representation jointly trained with 2D and 3D supervision, compressing TSDF volumes into a compact space. This enables scaling up to $64^{3}$ resolution completion while reducing GPU memory usage by $30\%$ compared to state-of-the-art [9, 7].
  • We introduce a guidance mechanism that combines cross-attention with ControlNet-inspired conditioning, enabling flexible integration of complementary modalities. We introduce a training strategy that simulates realistic partial shapes directly from 3D datasets, better reflecting real-world occlusions and noise while not making any assumptions about the input’s structure. This multimodal approach enables class-agnostic generation while producing more realistic, diverse, and high-fidelity completions compared to prior methods.
  • We achieve state-of-the-art performance on both synthetic and real-world data, outperforming class-specific models by $12\%$ and class-agnostic models by $47\%$ in $l_{1}$ reconstruction error.
@@ 2 Related Work
Traditional approaches to 3D shape completion primarily focus on repairing minor imperfections and geometrical gaps, employing strategies like Laplacian hole filling [20] and Poisson surface reconstruction [5]. While effective for filling small gaps in 3D shapes, these methods are of limited use in scenarios with larger gaps in the input scans. However, the advent of extensive 3D datasets has led to the development of learning-based approaches. While retrieval-based methods [4, 44, 3] extract the most fitting shapes from a database to complete the incomplete input, learning-based fitting methods [43, 13, 15, 16, 41] that aim to minimize the disparity between network-predicted shapes and actual ground truth have become more prevalent. For example, 3D-EPN [13] and Scan2Mesh [11] have employed 3D encoder-decoder architectures for predicting complete shapes from partial volumetric data. More recently, PatchComplete [32] introduces a novel design by learning multi-resolution patch priors, enhancing the completion of shapes, particularly for previously unseen categories. While these models can provide high-accuracy completions, their deterministic nature may lead to overfitting. Furthermore, they often lack the ability to generate finer details in shapes.
@@
Generative methods, including Generative Adversarial Networks (GANs) [45] and AutoEncoders [25], have recently been introduced to the field. These methods excel in generating a diverse range of plausible shapes from partial inputs. However, this often comes at the cost of accuracy, especially when a precise ground truth is available for training. Diffusion models have emerged as a popular family of generative models, renowned for their impressive quality, diversity, and expressiveness in tasks such as image synthesis [35], image editing [23], and text-to-image synthesis [31]. While extensively researched for 2D data, their application to 3D data has only been explored very recently, using different 3D representations such as point clouds [22, 48], neural radiance fields [24], or signed distance functions [8, 21]. SDFusion [7], specifically, introduces a latent diffusion model for shape generation aimed at mitigating the computational challenges associated with 3D representations, and conditions it on various modalities like images, text, and shapes. However, SDFusion frames shape completion as filling simple, artificially removed cubic regions, rather than reconstructing the irregular, realistic missing parts encountered in real-world scans. In contrast, our approach does not make any assumption on the partial shape as real-world scans are characterized by noise and varying degrees of incompleteness. Another approach, DiffComplete [9], offers a more realistic shape completion method but directly works on the 3D space and is therefore restricted to lower-resolution voxel grids of size $32^{3}$. As opposed to DiffComplete, our approach can achieve much higher resolution by encoding the 3D space to compact latent codes.
More recently, object completion has also been explored in a zero-shot setting using foundation models by using score distillation sampling [19]. However, these approaches require a lengthy per-object optimization process during test-time and rely on sparse yet mostly complete projections from a single viewpoint [1, 19]. Similarly, Gao et al. [17] propose multi-view completion using a diffusion model conditioned on multiple images. While promising, both directions make strong assumptions about the input data and incur substantial computational and memory costs. In contrast, our model makes no assumptions on the structure of its conditioning signal and remains comparatively lightweight in memory usage. Furthermore, most previous methods [32, 26, 24] work on the assumption that the object class is known prior to the completion, which makes them hard to deploy in real-world applications. In contrast, our model is class-agnostic, and outputs high-detailed, realistic completions, even for unseen classes.
Figure 2: We train a Vector Quantized Variational Autoencoder (VQ-VAE) to encode complete TSDF voxel grids $\mathbf{X}$ into a compact latent variable $\mathbf{z}$. Then, we use this learned latent space to train our conditional shape completion diffusion model. We corrupt the clean latent code $\mathbf{z}$ through a forward diffusion process $q(\mathbf{z}_{t}|\mathbf{z}_{-1})$ by gradually adding Gaussian noise until reaching the normally distributed variable $\mathbf{z}_{T}$. Our diffusion model is then trained to learn the reverse diffusion process $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})$ conditioned on a partial TSDF scan c of the shape by aggregating features from a control branch and/or a partial RGB image $\mathbf{I}$ of the shape by applying cross-attention on CLIP features of the given images. This process allows us to recover the latent code $\mathbf{\hat{z}}$, and finally, using the VQ-VAE’s decoder, the completed shape $\mathbf{\hat{X}}$.
@@ 3 Methodology
We approach 3D shape completion as a generative task based on a diffusion probabilistic model, aiming to produce a complete shape $\hat{\mathbf{X}}$ from a given partial scan represented by an RGB Image $\mathbf{I}$ and a partial 3D TSDF scan $\mathbf{c}$. While diffusion models have proven highly effective in generating high-resolution images, their direct application to high-resolution 3D shapes is hindered by the substantial demands on computation and memory. Consequently, 3D diffusion-based approaches are often confined to low-resolution voxel grids, typically around $32^{3}$  [7, 25]. To address this limitation, we first compress the 3D shape into a compact latent space using a Vector Quantized Variational Autoencoder (VQ-VAE), as detailed in Sec. 3.1. This VQ-VAE is trained with supervision from 3D TSDF as well as 2D depth and normal renderings from a raycasting-based volume rendering approach [16]. This compression strategy allows us to implement our diffusion model in a lower-dimensional space, a process referred to as latent diffusion (Sec. 3.2). Our model includes two distinct conditioning mechanisms (Sec. 3.3) in the generative process: the first leverages RGB Image $\mathbf{I}$’s CLIP features [30] via cross attention, and the second uses spatially localized conditions [46, 9] derived from the partial 3D TSDF scan $\mathbf{c}$. An overview of our approach is provided in Fig. 2.
@@
We employ a 3D variant of the VQ-VAE [40, 33] to condense the latent representations of 3D shapes. The architecture of the VQ-VAE consists of an encoder $\mathcal{E}$, a decoder $\mathcal{D}$, and a codebook $\bm{\mathcal{Z}}\in\mathbb{R}^{K_{\mathcal{Z}}\times D}$, where $K_{\mathcal{Z}}$ denotes the codebook size and $D$ the dimensionality of each latent vector. Compared to continuous latent spaces, discrete representations offer two major advantages in our setting. First, following prior work on latent diffusion models [34], vector quantization prevents the diffusion process from drifting into regions corresponding to invalid or implausible 3D shapes. Unlike VAEs, which tend to over-regularize and over-smooth latent codes, VQ-VAEs preserves the diversity of distinctly different shape categories, allowing the model to represent sharp structural variations more faithfully. Second, VQ-VAE substantially reduces memory usage during training, since only discrete codebook indices need to be stored instead of full-precision latent feature volumes. This two-stage compression makes high-resolution volumetric completion feasible, enabling us to scale to $64^{3}$ voxel grids with significantly lower memory requirements than continuous alternatives.
@@
Alongside 3D losses, our methodology incorporates 2D supervision from rendering the reconstructed TSDF grids $\mathbf{X^{\prime}}$, similar to SPSG [16]. We render both depth ($\mathbf{D}_{v}^{\prime}$) and normal images ($\mathbf{N}_{v}^{\prime}$) from $\mathbf{X^{\prime}}$ at each camera view $v$ in a differentiable manner via raycasting-based volume rendering. The 2D supervision losses employed for VQ-VAE training include the reconstruction loss and adversarial loss. The 2D reconstruction loss is an $l_{1}$ loss $\mathcal{L}_{R_{2D}}$ guiding depth and normal reconstruction:
@@
where $\mathbf{D}_{v}$ and $\mathbf{N}_{v}$ correspond to the ground-truth depth and normal values, and $\mathbf{D}_{v}^{\prime}$ and $\mathbf{N}_{v}^{\prime}$ are the rendered depth and normal images. In addition, we apply a patch-based adversarial loss $L_{A_{2D}}$ on normals $\mathbf{N}_{v}^{\prime}$ for realistic geometry reconstruction:
@@
where $\mathrm{Disc}(\cdot)$ denotes the CNN-based discriminator network. The final VQ-VAE objective $\mathcal{L}_{VQVAE}$ combines both 3D and 2D losses:
$\mathcal{L}_{VQVAE}=\mathcal{L}_{R}+\lambda\mathcal{L}_{C}+(1-\lambda)\mathcal{L}_{CB}+\gamma_{R}\mathcal{L}_{R_{2D}}+\gamma_{A}L_{A_{2D}},$ (3)